OpenStudy (anonymous):

Solve the system of equations and choose the correct ordered pair. 7x + 8y = 4 3x + 2y = 6 A. (4, -2) B. (2, 4) C. (2, 3) D. (4, -3)

OpenStudy (anonymous):

Can you use matrices?

OpenStudy (anonymous):

nope

OpenStudy (radar):

Since no instructions were given on which method, I would choose the "elimination" method. Start by multiplying the second equation by a -4 getting: -12x -8y =-24 now add the first equation. 7x +8y= 4 ---------------- -5x =-20 x= 4 after dividing by -5. 3(-4) + 2y =6 substituting x value in 2nd equation. -12 +2y = 6 2y =18 y=9

OpenStudy (radar):

Did you follow with understanding?

OpenStudy (anonymous):

ts wrong in a way thanks thou

OpenStudy (radar):

To check for accuracy try in original equation: 7(-4) + 8(9) =4 I can see problem, something amiss here. -28 +72=4 44=4 No way. Do you see where the error is located?

OpenStudy (radar):

Ah so I found it, I substituted a -4 rather than a +4 it should be: 3(4)+2y=6 12 +2y = 6 2y=6-12=-6 y=-3 so the ordered pair is (4, -3)